# Dance to Demise – How Massive Stars May Form Dense Circumstellar Shells Before Explosion

### v2

@@ 3.4 Dust driven wind
Beyond a few stellar radii, the temperature in RSG atmospheres drops below K, allowing the condensation of silicate and alumina grains. Radiation pressure on these dust grains can then drive a wind, especially when aided by prior pulsational levitation of gas (R. D. Gehrz & N. J. Woolf, 1971). Dust-driven winds are supported by observations of IR excesses and extended envelopes in nearby RSGs such as Betelgeuse (A. K. Dupree et al., 2022) and VY CMa (E. D. Beck et al., 2016). These winds often exhibit outflow velocities of 10–30 km s-1 and mass loss rates between and M⊙ (J. T. van Loon et al., 2005).
@@ 3.5 Enhanced mass loss due to binary companion
The majority of massive stars are believed to have experienced at least one episode of binary interaction in course of their evolution (N. Langer, 2012; H. Sana et al., 2012), although for RSGs, the binary fraction is estimated to be much lower both from observations (M. Dai et al., 2025a, b) and population synthesis predictions using the Binary Population and Spectral Synthesis code (BPASS; J. J. Eldridge et al. 2017; E. R. Stanway & J. J. Eldridge 2018; H. Stevance et al. 2020). Irradiation effects due to heating or the ram pressure from the wind of a stellar companion or even an explosion, as well as tidal forces exerted by a binary companion (either a star, planet or a compact remnant) causing redistribution of the angular momentum inside the star, can trigger enhanced mass loss.
Stars in binary systems may also accrete matter causing spin-up to critical rotation, thereby becoming partially unbound, resulting in mass loss through the inner Lagrangian point or even merging into a common envelope (M. Clayton et al., 2017; H. F. Stevance & J. J. Eldridge, 2021). Such evolutionary scenarios are also possible in X-ray binaries where the compact object (either a neutron star or black hole) accretes matter from its giant companion, making the latter lose mass through Roche lobe overflow, which could prevent it from evolving into a RSG and instead explode as a Type I b/c SN, a significant fraction of which are understood to originate from such envelope stripping by a binary companion (A. Grichener, 2023; T. Shenar, 2024).
@@ 4 RSG models with pulsations
We model massive stars with initial masses between M⊙ which evolve into red supergiants (RSGs) using the open-source stellar evolution code Modules for Experiments in Stellar Astrophysics (MESA; B. Paxton et al., 2011, 2013, 2015, 2018, 2019) - version 24.08.1. These models are computed hydrostatically from the pre-main sequence through to the end of core helium burning, and thereon hydrodynamically until core oxygen exhaustion. We use a smaller nuclear network, approx21_cr60_plus_co56.net (F. X. Timmes, 1999) until core carbon depletion, followed by a larger network, mesa_128.net (F. X. Timmes, 1999; F. X. Timmes et al., 2000) to better capture the energy generation in advanced phases of nuclear burning, due to limited computational resources. Our models also incorporate time-dependent convection as implemented in MESA by A. S. Jermyn et al. (2023) to better approximate the turbulent conditions within the convective envelopes of RSGs, than it is possible using the standard mixing length approach. We use MESA’s inbuilt Dutch wind scheme until the end of core helium burning, scaled by a factor dutch_scaling_factor, which gives RSG mass loss rates consistent with observations of large samples of RSGs in the Milky Way, the Magellanic Clouds (M. Yang et al., 2023; K. Antoniadis et al., 2024; J. Wen et al., 2024) as well as M31, M33 (P. Massey et al., 2023) and other nearby star clusters with coeval populations of RSGs from which mass-dependent mass loss rates have been inferred (E. R. Beasor et al., 2023; L. Decin et al., 2024). Beyond core helium exhaustion, we switch to our custom mass loss prescription as described below in 4.1 to model the effect of radial pulsations on mass loss in RSGs (). We also investigate late mass ejections in the subsequent evolution of our RSG models before CC, due to dynamical instabilities post-shock breakout in the pulsating stellar envelope, as described in 4.2. In 4.3 we describe the model for line-driven acceleration of the RSG wind generated through the various mass loss mechanisms mentioned above. Finally in 4.3.2 we calculate the CSM density profiles and the column density along line-of-sight assuming spherically symmetric wind using the accelerated velocity profile for comparison with few well-known Type II SNe (Section 5).
@@ 4.1 Pulsation-Driven Superwind (PDSW)
Red Supergiants (RSGs) often exhibit long-period variability (M. D. Soraisam et al., 2018; Y. Ren et al., 2019; B. Hsu et al., 2024) that has been associated with changes in stellar radii due to excitation of different modes of pulsational instabilities through the -mechanism (H. Saio et al., 2016). Theoretically, A. Heger et al. (1997) demonstrated the growth of such pulsational instabilities in hydrodynamical evolutionary calculations with sufficient temporal resolution and found that the growth rates in their non-linear numerical calculations to be proportional to the luminosity () to mass () ratio. S.-C. Yoon & M. Cantiello (2010) showed that the growth rate () of the amplitude of the surface velocities increases with ratio and/or decreasing thermal (Kelvin-Helmholtz) time scale of the envelope (), according to the relation:
@@
where is the stellar radius, ( is the mass of the H-rich envelope ) and which are calibrated against their hydrodynamic models evolved until shock breakout at the surface. Figure 1 shows a linear least-squares fit for the relation given by Equation 1 for hydrodynamic MESA model sequences with initial masses between M⊙ evolved with timesteps until shocks develop at the stellar surface, which gives , in close agreement with the values of S.-C. Yoon & M. Cantiello (2010).
@@
The top panel of Figure 2 shows a typical hydrodynamical model for a MESA run with initial mass M⊙ illustrating the growth of the radial pulsations until shock break-out at the surface where the Mach number reaches unity. The bottom panel of Figure 2 shows the corresponding evolution of a set of MESA runs with varying initial masses between M⊙ in the Hertzspring-Russell (HR) diagram. Capturing these radial pulsations with the typical time period, , requires using much smaller time steps (B. Paxton et al., 2013), that makes modeling such pulsations a challenging task for evolutionary codes which typically use much larger time steps during the onset and growth of these pulsations (B. Paxton et al., 2013; V. A. Bronner et al., 2025). We do not attempt to resolve these pulsations for our full evolutionary runs (from pre main-sequence until core neon/oxygen depletion) for computational feasibility but incorporate potential impact of the growth of such pulsations observed in advanced phases (beyond core-helium burning) on the mass loss history of our stellar models in a phenomenological approach, similar to that of S.-C. Yoon & M. Cantiello (2010). To do this, we evolve a set of hydrostatic MESA models until end of core-helium burning using the standard Dutch wind scheme in MESA, scaled down by a factor of for reasons explained in Section 4. Thereafter, each run is continued with the hydrodynamic term on and replacing the wind mass loss rates with the ”superwind” (SW) mass loss prescription of S.-C. Yoon & M. Cantiello (2010) given by
@@
for and , using the growth rate expression as given in Equation 1 with our calibrated constants. Here, is the mass loss rate from de C. de Jager et al. (1988) scaled down by a factor of , and is a free parameter which sets the strength and consequently the duration of the SW-phase (Section 4.2). Unlike S.-C. Yoon & M. Cantiello (2010) who used a fixed , we conduct a thorough exploration of the dependence of the SW on both the initial mass M⊙ and value of , motivated by observational evidence for enhanced mass loss in late stage of evolution of RSG progenitors of Type II SNe.
@@
Figure 3 illustrates a sequence of such MESA models for M⊙, showing the evolution of the mass loss rate as a function of the logarithm of the time to CC, measured from right to left. Thus, the zero of the x-axis represents one year from CC, approximately corresponding to the duration of oxygen-burning in our models, which is the final phase up to which we have evolved all our model sequences. It is readily seen that the model with evolves with an enhanced mass loss rate over the last years, which is approximately two orders of magnitude higher compared to the model sequences evolved with the Dutch prescription in MESA, scaled down by a factor of , obtained by setting in Equation 2. At lower values of either the initial mass or , mass loss is less enhanced compared to the corresponding Dutch wind values but the qualitative picture is similar for lower mass progenitors with similar or lower values of (see Figure 5(a)).
@@ 4.2 Dynamical ejections post shock break-out
The SW prescription of S.-C. Yoon & M. Cantiello (2010) is strictly applicable only to the exponential growth phase of radial pulsations, which subsequently get damped by shocks that develop within the RSG envelope where the radial velocity of the loosely bound layers becomes supersonic. M. Clayton (2018) investigated the problem of dynamical stability of RSG envelopes in the shock-dominated regime, in terms of the pressure-weighted, volume averaged first adiabatic exponent (see Section 3.4 in M. Clayton et al., 2017) whose value less than at any point within the stellar envelope is indicative of instability to dynamical ejections from that point outwards (A. Lobel, 2001) - commonly occurring in regions above the hydrogen and helium ionization zones. Detailed shock-resolving hydrodynamic MESA models of M. Clayton (2018) showed launch of ejections upon the surface breakout of compression shocks followed by sufficiently fast expansion that raises material on to an escape trajectory. M. Clayton (2018) implemented a wind-mass loss scheme in order to excise any such contiguous unbound sub-surface layer of mass (with radial velocity exceeding the local escape velocity), by applying a mass loss rate of corresponding to a time-scale which is at least one order of magnitude below both the dynamical and thermal time-scales of the star. A subset of their extensive suite of hydrodynamic simulations of such shock-driven mass loss resulted in episodes of repeated mass ejection events lasting for upto -s of years. Almost all cases that converged upon a solution after showing such episodes of repeated mass ejections had , for which M. Clayton (2018) extracted time-averaged mass loss rates to propose the following mass loss prescription:
@@
above a cut-off value of between . Hence, we employ this mass loss recipe to model the possibility of such repeating mass ejections occurring post shock break-out in pulsationally unstableRSG envelopes when the luminosity to mass ratio of the star exceeds the cut-off value of , which is on the conservative side, given the uncertainties in determining the exact value of this cut-off (M. Clayton, 2018). We do not consider any other mass loss mechanism in this regime, so as to avoid duplicating effects of any similar mechanism(s) triggering mass loss in the final years leading to CC. Furthermore, the above recipe does not account for other dynamically unstable scenarios which lead to single ejection episodes, also identified in the simulations of M. Clayton (2018), that occur on a much shorter timescale to be accounted for in our evolutionary calculations.
@@
Figure 4 shows such a model sequence of initial mass M⊙ evolved with both the prescriptions given by Equations 2 (for ) and 3 implemented in MESA, with a transition at . It can be seen that the onset of the dynamical ejections marks a significant enhancement in the mass loss rates (around an order of magnitude) compared to the superwind phase, which has important implications for the CSM formed by these winds close to the star, as in this model the ejections occur less than years from CC. Figure 5 shows a set of evolutionary calculations using both the mass loss prescriptions for PDSW (Equation 2) and subsequent dynamical ejections (Equation 3). From our grid of MESA models, we find that only model sequences with initial mass M⊙ evolve to enter the regime for which dynamical ejections are expected (as described in 4.2) for . Furthermore, depending on the initial mass, there is a cut-off value of above which the dynamical ejections lead to elevated mass loss ( M⊙ ).
@@ 4.3.1 Acceleration of the RSG wind
Following T. J. Moriya et al. (2018), we consider acceleration of the stellar wind – which starts with a velocity at the launch radius , which is is assumed to be just the stellar radius – out to large radii where the wind velocity gradually reaches a constant terminal speed, . Since the exact mechanism of acceleration of the wind remains unknown, the kinematics is usually parametrized in form of the following simple -law for the wind velocity (H. J. G. L. M. Lamers & J. P. Cassinelli, 1999):
where the value of determines the shape of the wind profile. Usually, both and are obtained through spectral line fitting for which has been used in the literature (R.-P. Kudritzki & J. Puls, 2000). RSGs known to experience slower wind acceleration as compared to hotter (OB) stars i.e. (R. Baade et al., 1996; P. D. Bennett, 2010). For the following comparisons of our results with CSM properties of observed SNe (Section 5), we adopt a constant value of , consistent with hydrodynamical simulations (M. Curé & I. Araya, 2023) as well as observations of single RSGs (P. D. Bennett 2010; see Appendix A) while the values of the terminal velocity, , are used specific to each SN as inferred from fitting of spectral lines.
@@ 5.1 SN 2023ixf
SN 2023ixf is one of the nearest Type II SNe observed in recent years, with extensive follow-up providing stringent constraints on the structure and density of its CSM (W. V. Jacobson-Galán, 2025). Flash-ionization signatures in its spectra indicate that its red supergiant (RSG) progenitor underwent enhanced mass loss ( M⊙ ) during the final decade before CC, producing a compact dense shell (CDS) around it (J. E. Jencson et al., 2023; W. V. Jacobson-Galán et al., 2023; E. A. Zimmerman et al., 2024; N. Smith et al., 2023; K. A. Bostroem et al., 2024). Multi-epoch NuSTAR, XMM-Newton, Swift-XRT and Chandra X-ray observations (P. Chandra et al., 2024) as well as radio follow-up with the NOEM, VLA, GMRT (N. A. J. et al., 2025) infer mass loss rates of M⊙ at around cm, indicating a variable mass loss rates over the last yrs of the progenitor’s life. To model this scenario, in Figure 6(a), we show our MESA model sequence for a progenitor of initial mass M⊙ with a superwind parameter . This model enters the regime of dynamical ejections yrs prior to CC. Figure 6(a) shows the mass loss history of this model sequence, which shows the superwind phase starting at yrs before CC, with M⊙ , followed by episodes of dynamical ejections with averaged mass loss rates exceeding M⊙ in the final decade before explosion.
@@
Using a terminal wind speed of measured from He I emission (D. Dickinson et al., 2025; W. V. Jacobson-Galán et al., 2025), we compare our model predictions with inferred CSM densities from multi-epoch observations across optical, X-ray and radio wavelengths (N. A. J. et al., 2025), as shown in Figure 6(b). We obtain good agreement with both the early and late epoch observations, using the mass loss history of our M⊙, model sequence that leads to high CSM densities () close to the progenitor surface ( cm).
@@
In Figure 6(c), we compare the column density of neutral hydrogen, computed using Equation (7), with measured X-ray values as reported in P. Chandra et al. (2024) and N. A. J. et al. (2025), for which we use the evolution of the shock radius with time as per the estimates of P. Chandra et al. (2024), to convert the epochs into . We observe that, in general, we overpredict the column densities at most epochs where we have observed X-ray measurements. This could arise due to possible global asymmetries in the CSM, as hinted by the rapid spectropolarimetric properties at early times in the ejecta (S. S. Vasylyev et al., 2023; M. Shrestha et al., 2025) and/or from clumps of material along the line of sight, as indicated by the detection of Fe emission in the X-ray spectra (N. A. J. et al., 2025). While such effects have been attributed to the ejecta geometry (L. Ferrari et al., 2024) and/or binary interaction (P. D. Michel et al., 2025), regardless of the exact cause of their origins, either scenario would result in underestimation of the integrated column densities along the line of sight in observations, which would make them consistent with our systematically higher model estimates.

### v3

@@ 3.4 Dust driven wind
Beyond a few stellar radii, the temperature in RSG atmospheres drops below K, allowing the condensation of silicate and alumina grains. Radiation pressure on these dust grains can then drive a wind, especially when aided by prior pulsational levitation of gas (1971ApJ...165..285G). Dust-driven winds are supported by observations of IR excesses and extended envelopes in nearby RSGs such as Betelgeuse (Dupree_2022) and VY CMa (DeBeck2016ALMAOO). These winds often exhibit outflow velocities of 10–30 km s-1 and mass loss rates between and M⊙ (2005A&A...438..273V).
@@ 3.5 Enhanced mass loss due to binary companion
The majority of massive stars are believed to have experienced at least one episode of binary interaction in course of their evolution (2012ARA&A..50..107L; 2012Sci...337..444S), although for RSGs, the binary fraction is estimated to be much lower both from observations (2025MNRAS.539.1220D; 2025ApJ...988...60D) and population synthesis predictions using the Binary Population and Spectral Synthesis code (BPASS; 2017PASA...34...58E; 2018MNRAS.479...75S; 2020JOSS....5.1987S). Irradiation effects due to heating or the ram pressure from the wind of a stellar companion or even an explosion, as well as tidal forces exerted by a binary companion (either a star, planet or a compact remnant) causing redistribution of the angular momentum inside the star, can trigger enhanced mass loss.
Stars in binary systems may also accrete matter causing spin-up to critical rotation, thereby becoming partially unbound, resulting in mass loss through the inner Lagrangian point or even merging into a common envelope (2017MNRAS.470.1788C; 2021MNRAS.504L..51S). Such evolutionary scenarios are also possible in X-ray binaries where the compact object (either a neutron star or black hole) accretes matter from its giant companion, making the latter lose mass through Roche lobe overflow, which could prevent it from evolving into a RSG and instead explode as a Type I b/c SN, a significant fraction of which are understood to originate from such envelope stripping by a binary companion (2023MNRAS.523..221G; 2024IAUS..361..465S).
@@ 4 RSG models with pulsations
We model massive stars with initial masses between M⊙ which evolve into red supergiants (RSGs) using the open-source stellar evolution code Modules for Experiments in Stellar Astrophysics (MESA; Paxton2011; Paxton2013; Paxton2015; Paxton2018; Paxton2019) - version 24.08.1. These models are computed hydrostatically from the pre-main sequence through to the end of core helium burning, and thereon hydrodynamically until core oxygen exhaustion (for the full MESA implementation of our setup, refer to the Zenodo repository: sengupta_2025_17605468). We use a smaller nuclear network, approx21_cr60_plus_co56.net (1999ApJS..124..241T) until core carbon depletion, followed by a larger network, mesa_128.net (1999ApJS..124..241T; Timmes_2000) to better capture the energy generation in advanced phases of nuclear burning, due to limited computational resources. Our models also incorporate time-dependent convection as implemented in MESA by 2023ApJS..265...15J to better approximate the turbulent conditions within the convective envelopes of RSGs, than it is possible using the standard mixing length approach. We use MESA’s inbuilt Dutch wind scheme until the end of core helium burning, scaled by a factor dutch_scaling_factor, which gives RSG mass loss rates consistent with observations of large samples of RSGs in the Milky Way, the Magellanic Clouds (2023A&A...676A..84Y; antoniadis_2024; 2024AJ....167...51W) as well as M31, M33 (2023ApJ...942...69M) and other nearby star clusters with coeval populations of RSGs from which mass-dependent mass loss rates have been inferred (10.1093/mnras/stad1818; 2024A&A...681A..17D). Beyond core helium exhaustion, we switch to our custom mass loss prescription as described below in 4.1 to model the effect of radial pulsations on mass loss in RSGs (). We also investigate late mass ejections in the subsequent evolution of our RSG models before CC, due to dynamical instabilities post-shock breakout in the pulsating stellar envelope, as described in 4.2. In 4.3 we describe the model for line-driven acceleration of the RSG wind generated through the various mass loss mechanisms mentioned above. Finally in 4.3.2 we calculate the CSM density profiles and the column density along line-of-sight assuming spherically symmetric wind using the accelerated velocity profile for comparison with few well-known Type II SNe (Section 5).
@@ 4.1 Pulsation-Driven Superwind (PDSW)
Red Supergiants (RSGs) often exhibit long-period variability (2018ApJ...859...73S; 2019ApJS..241...35R; 2024arXiv240807874H) that has been associated with changes in stellar radii due to excitation of different modes of pulsational instabilities through the -mechanism (2016IAUFM..29B.573S). Theoretically, 1997A&A...327..224H demonstrated the growth of such pulsational instabilities in hydrodynamical evolutionary calculations with sufficient temporal resolution and found that the growth rates in their non-linear numerical calculations to be proportional to the luminosity () to mass () ratio. 2010ApJ...717L..62Y showed that the growth rate () of the amplitude of the surface velocities increases with ratio and/or decreasing thermal (Kelvin-Helmholtz) time scale of the envelope (), according to the relation:
@@
where is the stellar radius, ( is the mass of the H-rich envelope ) and which are calibrated against their hydrodynamic models evolved until shock breakout at the surface. Figure 1 shows a linear least-squares fit for the relation given by Equation 1 for hydrodynamic MESA model sequences with initial masses between M⊙ evolved with timesteps until shocks develop at the stellar surface, which gives , in close agreement with the values of 2010ApJ...717L..62Y.
@@
The top panel of Figure 2 shows a typical hydrodynamical model for a MESA run with initial mass M⊙ illustrating the growth of the radial pulsations until shock break-out at the surface where the Mach number reaches unity. The bottom panel of Figure 2 shows the corresponding evolution of a set of MESA runs with varying initial masses between M⊙ in the Hertzspring-Russell (HR) diagram. Capturing these radial pulsations with the typical time period, , requires using much smaller time steps (Paxton2013), that makes modeling such pulsations a challenging task for evolutionary codes which typically use much larger time steps during the onset and growth of these pulsations (Paxton2013; 2025A&A...703A..61B). We do not attempt to resolve these pulsations for our full evolutionary runs (from pre main-sequence until core neon/oxygen depletion) for computational feasibility but incorporate potential impact of the growth of such pulsations observed in advanced phases (beyond core-helium burning) on the mass loss history of our stellar models in a phenomenological approach, similar to that of 2010ApJ...717L..62Y. To do this, we evolve a set of hydrostatic MESA models until end of core-helium burning using the standard Dutch wind scheme in MESA, scaled down by a factor of for reasons explained in Section 4. Thereafter, each run is continued with the hydrodynamic term on and replacing the wind mass loss rates with the ”superwind” (SW) mass loss prescription of 2010ApJ...717L..62Y given by
@@
for and , using the growth rate expression as given in Equation 1 with our calibrated constants. Here, is the mass loss rate from de 1988A&AS...72..259D scaled down by a factor of , and is a free parameter which sets the strength and consequently the duration of the SW-phase (Section 4.2). Unlike 2010ApJ...717L..62Y who used a fixed , we conduct a thorough exploration of the dependence of the SW on both the initial mass M⊙ and value of , motivated by observational evidence for enhanced mass loss in late stage of evolution of RSG progenitors of Type II SNe.
@@
Figure 3 illustrates a sequence of such MESA models for M⊙, showing the evolution of the mass loss rate as a function of the logarithm of the time to CC, measured from right to left. Thus, the zero of the x-axis represents one year from CC, approximately corresponding to the duration of oxygen-burning in our models, which is the final phase up to which we have evolved all our model sequences. It is readily seen that the model with evolves with an enhanced mass loss rate over the last years, which is approximately two orders of magnitude higher compared to the model sequences evolved with the Dutch prescription in MESA, scaled down by a factor of , obtained by setting in Equation 2. At lower values of either the initial mass or , mass loss is less enhanced compared to the corresponding Dutch wind values but the qualitative picture is similar for lower mass progenitors with similar or lower values of (see Figure 5).
@@ 4.2 Dynamical ejections post shock break-out
The SW prescription of 2010ApJ...717L..62Y is strictly applicable only to the exponential growth phase of radial pulsations, which subsequently get damped by shocks that develop within the RSG envelope where the radial velocity of the loosely bound layers becomes supersonic. clayton2018a investigated the problem of dynamical stability of RSG envelopes in the shock-dominated regime, in terms of the pressure-weighted, volume averaged first adiabatic exponent (see Section 3.4 in 2017MNRAS.470.1788C) whose value less than at any point within the stellar envelope is indicative of instability to dynamical ejections from that point outwards (2001ApJ...558..780L) - commonly occurring in regions above the hydrogen and helium ionization zones. Detailed shock-resolving hydrodynamic MESA models of clayton2018a showed launch of ejections upon the surface breakout of compression shocks followed by sufficiently fast expansion that raises material on to an escape trajectory. clayton2018a implemented a wind-mass loss scheme in order to excise any such contiguous unbound sub-surface layer of mass (with radial velocity exceeding the local escape velocity), by applying a mass loss rate of corresponding to a time-scale which is at least one order of magnitude below both the dynamical and thermal time-scales of the star. A subset of their extensive suite of hydrodynamic simulations of such shock-driven mass loss resulted in episodes of repeated mass ejection events lasting for upto -s of years. Almost all cases that converged upon a solution after showing such episodes of repeated mass ejections had , for which clayton2018a extracted time-averaged mass loss rates to propose the following mass loss prescription:
@@
above a cut-off value of between . Hence, we employ this mass loss recipe to model the possibility of such repeating mass ejections occurring post shock break-out in pulsationally unstable RSG envelopes when the luminosity to mass ratio of the star exceeds the cut-off value of , which is on the conservative side, given the uncertainties in determining the exact value of this cut-off (clayton2018a). We do not consider any other mass loss mechanism in this regime, so as to avoid duplicating effects of any similar mechanism(s) triggering mass loss in the final years leading to CC. Furthermore, the above recipe does not account for other dynamically unstable scenarios which lead to single ejection episodes, also identified in the simulations of clayton2018a, that occur on a much shorter timescale to be accounted for in our evolutionary calculations.
@@
Figure 4 shows such a model sequence of initial mass M⊙ evolved with both the prescriptions given by Equations 2 (for ) and 3 implemented in MESA, with a transition at . It can be seen that the onset of the dynamical ejections marks a significant enhancement in the mass loss rates (around an order of magnitude) compared to the superwind phase, which has important implications for the CSM formed by these winds close to the star, as in this model the ejections occur less than years from CC.
Figure 5 shows a set of evolutionary calculations using both the mass loss prescriptions for PDSW (Equation 2) and subsequent dynamical ejections (Equation 3). From our grid of MESA models, we find that only model sequences with initial mass M⊙ evolve to enter the regime for which dynamical ejections are expected (as described in 4.2) for . Furthermore, depending on the initial mass, there is a cut-off value of above which the dynamical ejections lead to elevated mass loss ( M⊙ ).
@@ 4.3.1 Acceleration of the RSG wind
Following 2018MNRAS.476.2840M, we consider acceleration of the stellar wind – which starts with a velocity at the launch radius , which is is assumed to be just the stellar radius – out to large radii where the wind velocity gradually reaches a constant terminal speed, . Since the exact mechanism of acceleration of the wind remains unknown, the kinematics is usually parametrized in form of the following simple -law for the wind velocity (1999isw..book.....L):
where the value of determines the shape of the wind profile. Usually, both and are obtained through spectral line fitting for which has been used in the literature (2000ARA&A..38..613K). RSGs known to experience slower wind acceleration as compared to hotter (OB) stars i.e. (1996ApJ...466..979B; 2010ASPC..425..181B). For the following comparisons of our results with CSM properties of observed SNe (Section 5), we adopt a constant value of , consistent with hydrodynamical simulations (galaxies11030068) as well as observations of single RSGs (2010ASPC..425..181B; see Figure 11 in Appendix A) while the values of the terminal velocity, , are used specific to each SN as inferred from fitting of spectral lines.
@@ 5.1 SN 2023ixf
SN 2023ixf is one of the nearest Type II SNe observed in recent years, with extensive follow-up providing stringent constraints on the structure and density of its CSM (2025arXiv250708078J). Flash-ionization signatures in its spectra indicate that its red supergiant (RSG) progenitor underwent enhanced mass loss ( M⊙ ) during the final decade before CC, producing a compact dense shell (CDS) around it (2023ApJ...952L..30J; 2023ApJ...954L..42J; 2024Natur.627..759Z; Smith2023ixf; 2024ApJ...973L..47B). Multi-epoch NuSTAR, XMM-Newton, Swift-XRT and Chandra X-ray observations (Chandra_2024) as well as radio follow-up with the NOEM, VLA, GMRT (2025ApJ...985...51A) infer mass loss rates of M⊙ at around cm, indicating a variable mass loss rates over the last yrs of the progenitor’s life. To model this scenario, in the top panel of Figure 6, we show our MESA model sequence for a progenitor of initial mass M⊙ with a superwind parameter . This model enters the regime of dynamical ejections yrs prior to CC. The top panel of Figure 6 shows the mass loss history of this model sequence, which shows the superwind phase starting at yrs before CC, with M⊙ , followed by episodes of dynamical ejections with averaged mass loss rates exceeding M⊙ in the final decade before explosion.
@@
Using a terminal wind speed of measured from He I emission (2025ApJ...984...71D; 2025arXiv250811747J), we compare our model predictions with inferred CSM densities from multi-epoch observations across optical, X-ray and radio wavelengths (2025ApJ...985...51A), as shown in the middle panel of Figure 6. We obtain good agreement with both the early and late epoch observations, using the mass loss history of our M⊙, model sequence that leads to high CSM densities () close to the progenitor surface ( cm).
@@
In the bottom panel of Figure 6, we compare the column density of neutral hydrogen, computed using Equation (7), with measured X-ray values as reported in Chandra_2024 and 2025ApJ...985...51A, for which we use the evolution of the shock radius with time as per the estimates of Chandra_2024, to convert the epochs into . We observe that, in general, we overpredict the column densities at most epochs where we have observed X-ray measurements. This could arise due to possible global asymmetries in the CSM (2025A&A...694A.319K), as hinted by the rapid spectropolarimetric properties at early times in the ejecta (2023ApJ...955L..37V; 2025ApJ...982L..32S) and/or from clumps of material along the line of sight, as indicated by the detection of Fe emission in the X-ray spectra (2025ApJ...985...51A). While such effects have been attributed to the ejecta geometry (2024A&A...687L..20F) and/or binary interaction (10.1093/mnras/staf443), regardless of the exact cause of their origins, either scenario would result in underestimation of the integrated column densities along the line of sight in observations, which would make them consistent with our systematically higher model estimates.